# HOIDiNi: Human-Object Interaction through Diffusion Noise Optimization

### v1

@@ 4.2. CPHOI Diffusion Model
CPHOI generates the next motion segment $s^{n}=[F_{i}]_{i=1}^{L}$ of $L$ frames, given the previous segment, $s^{n-1}$, the object geomtry, and a text prompt that describes the interaction. For each denoising step $t\in[0,T]$, the inputs to the transformer decoder backbone are the frames of the previous segment $s^{n-1}$, followed by the current segment to be denoised, $s^{n}_{t}$. The model predicts the clean version of the segment, $\hat{s}^{n}_{0}$.
The object’s geometry, the text condition, and the timestep $t$ are encoded into a sequence of latent embeddings and injected through the cross-attention of each transformer layer, at each denoising step. To encode the geometry of the interacted object, we use a shape encoder with a PointNet++ (Qi et al., 2017) architecture, which is trained simultaneously with the diffusion model. We sample a subset of the mesh vertices and encode them into a latent descriptor per vertex. The text prompt is encoded into a sequence of embeddings using a pre-trained and fixed DistilBERT (Sanh, 2019) model. The diffusion timestep $t$ is embedded using a standard positional embedding denoted $E_{t}$.
@@
4.3. Two-phase Generation Optimization
Accurate body-object contacts are a key component of plausible HOI motion, but represent a discontinuous space, with discrete and unstable decision making. Thus we found the straight forward optimization, using a single step challenging in practice. To address this, we separate the process into two phases, each solving a different part of the problem variables:
Phase 1: Object-Centric
In this phase, we optimize the object related part only to outline the motion, based on the given prompt, object geometry, and scene constraints (e.g., table surface). Although the full output of the model is the triplet $(CP,H,O)$—representing the contact-pair sequence, human motion, and object motion respectively—only $(CP)$ and $(O)$ are considered in this stage. We found weight sharing between phases to improve performance. The objective $\mathcal{L}_{Object}$ for this phase is defined to avoid object-scene penetrations and optionally to guide the object toward a set of target poses in specified frames:
@@ 4.3. Two-phase Generation Optimization
Figure 5. Contact Pairs. CPHOI predicts precise, semantically meaningful contact points between the hand and object. Each contact pair is visualized with matching colored spheres.
This objective comprises a penetration term between the object and the scene $\mathcal{L}_{P_{OS}}$, a goal term encouraging the object to achieve specified poses in specified times $\mathcal{L}_{Goal}$, and the static term, $\mathcal{L}_{Static}$, that encourage the object to stay static in frames without predicted contact.
Penetration Loss.
To measure interpenetration between two watertight meshes, denoted $M_{a}$ and $M_{b}$, we define a bidirectional penetration loss that penalizes vertices of one mesh that are located inside the volume enclosed by the other. Each mesh $M$ consists of a set of vertices $\mathcal{V}$ and a set of polygons $\mathcal{F}$. Each polygon $f\in\mathcal{F}$ is associated with an outward-facing normal vector $\mathbf{n}_{f}$.
Figure 6. Comparisons. HOIDiNi generates semantically correct and accurate interaction with the gaming controller. From left to right: IMoS (Ghosh et al., 2023) optimization yields inferior contacts and unrealistic motion; CPHOI inference only generates decent poses but fails to satisfy contacts; Our losses with Classifier Guidance brings the object insufficiently closer. Replacing our contact-point prediction with the popular nearest-neighbor heuristic fails to choose correct contacts, in contrast to our plausible and human-like result.
For each vertex $\mathbf{v}\in\mathcal{V}_{a}$, we identify whether it is inside the other mesh simply through randomly casting a single ray, and checking the normal direction of the intersected triangle if it exsists. we denote the vertices found as inside $\mathcal{V}_{a}^{\text{in}}\subset\mathcal{V}_{a}$.
@@
$\mathcal{L}_{\text{Penetration}}=\mathcal{L}_{\text{pen}}(M_{a}\rightarrow M_{b})+\mathcal{L}_{\text{pen}}(M_{b}\rightarrow M_{a})$
@@
Phase 2: Human-Centric Phase
This phase refines the human motion to conform to the fixed object motion and contact sequence $(CP,O)$ produced in Phase 1. The objective of this diffusion-based optimization is defined as:
representing human-object, human-scene, and human-human penetration losses, respectively. Specifically, $\mathcal{L}_{P_{HO}}$ penalizes interpenetration between the human and the object, $\mathcal{L}_{P_{HS}}$ prevents collisions with the static scene, and $\mathcal{L}_{P_{HH}}$ reduces self-intersections within the human mesh, particularly between the hands.
The contact loss $\mathcal{L}_{C}$ uses $L_{2}$ distance to push hand vertex locations to targets on the moving object, as dictated by $CP$ and $O$ from the previous phase. The penetration loss terms are similar to the penetration term $\mathcal{L}_{P_{OS}}$ defined in phase 1.
Together, these losses represent physical plausibility, encouraging surface contacts while avoiding penetration, and the optimization scheme ensures realism and human likeness.
GT 0.995 74.6% 0.194 5.0 $\pm$ 1.8 2.6 $\pm$ 1.8
Ours 0.159 0.996 0.121 62.3% 0.245 6.8 $\pm$ 2.9 2.3 $\pm$ 4.7
IMoS 0.205 1.026 0.121 46.5% 0.204 3.0 $\pm$ 7.7 52.2 $\pm$ 53.8
Single-Phase 0.221 1.026 0.140 43.0% 0.241 7.8 $\pm$ 3.6 20.1 $\pm$ 21.7
Higher Penetration Coefficient $\lambda_{P_{HO}}$ 0.156 1.015 0.144 57.0% 0.225 5.6 $\pm$ 2.9 5.9 $\pm$ 7.2
Nearest Neighbors instead of Contact Pairs 0.162 1.040 0.153 57.9% 0.245 7.7 $\pm$ 3.8 4.3 $\pm$ 5.0
No Jitter Loss 0.160 1.026 0.143 60.5% 0.220 7.1 $\pm$ 2.7 1.9 $\pm$ 4.0
Table 1. Quantitative Results and Ablation Study. Comparison on the GRAB dataset (Taheri et al., 2020a) using IMoS-defined (Ghosh et al., 2023) metrics for motion realism, along with average floating and penetration errors (in millimeters). $\rightarrow$ indicates that better is closer to ground-truth performance.
5. Experiments
Our code, model checkpoints, and additional qualitative results are available at: https://hoidini.github.io.
5.1. Evaluation Setting
Implementation Details.
CPHOI is implemented as a Transformer decoder architecture with 8 layers and a hidden dimension of 512. Our point-wise object embedding network is a PointNet++  (Qi et al., 2017) fed with 512 randomly sampled vertices from the conditioned object. We condition the motion on a prefix of 15 frames and generate 100 frames. The model is trained with DDPM (Ho et al., 2020a); DDIM (Song et al., 2020) is used at inference. Both training and inference use 8 diffusion steps. We train the model for 120K iterations using a batch size of 64. Each phase runs for 350 optimization steps with a learning rate of 0.9 and 20 warm-up steps. We use the Adam optimizer (Kingma, 2014) for the noise optimization procedure. To reduce memory we focus only on the MANO  (Romero et al., 2017b) subset of vertices of the SMPL-X human body and simplify it further to 1,100 vertices for each hand. The object meshes are simplified to 3,000 faces. At each phase, we apply a perturbation scale of $10^{-6}$ and a difference penalty of $10^{-6}$ from the initially sampled noise. We set the loss coefficients as follows: $\lambda_{C}=0.95,\ \lambda_{Foot}=0.5,\ \lambda_{Jitter}=0.00001,\ \lambda_{P_{HO}}=0.05,\ \lambda_{P_{HS}}=0.2,\ \lambda_{P_{HH}}=0.05,\ \lambda_{P_{OS}}=1.2,\ \lambda_{Goal}=0.5,\ \lambda_{Static}=0.9$.
Data.
We evaluate our approach using the GRAB dataset (Taheri et al., 2020b), which includes SMPL-X (Pavlakos et al., 2019b) human motion parameters, object 6DoF trajectories, and per-vertex contact force data for both human and object meshes. The dataset comprises 1,334 motion samples from 10 different subjects. Since different human subjects in the dataset exhibit varying shape parameters $\beta$, and considering the dataset’s relatively small scale compared to datasets like HumanML3D, we re-target all motions to a standardized, neutral SMPL-X model with shape parameters fixed to $\beta=0$.
To help combat the sparsity and minimal diversity of the dataset, We also condition the model on text descriptions. To capture both the object and the associated action, we use a language model (GPT-4o) to generate simple prompts based on each sample name, including the relevant action and object.
Baselines
We compare our approach against IMoS (Ghosh et al., 2023) and several internal baselines. One baseline applies our model at inference time without any optimization. Another replaces the diffusion noise optimization with classifier guidance. The third substitutes the predicted contact pairs with nearest-neighbor-based pairs. To enable a fair comparison with classifier guidance, we adopt DDPM hyper-sampling and increase the number of denoising steps from 8 to 500, consistent with Peng et al. (Peng et al., 2023). We report results using the optimal guidance scale. For the nearest-neighbor baseline, contact assignment is updated at each optimization step based on the current nearest neighbors. Additionally, we explore increasing the penetration loss weight by setting $\lambda_{P_{HO}}=0.9$.
Metrics
We evaluated our method across two dimensions: motion realism and interaction accuracy. Following the approach outlined in IMoS (Ghosh et al., 2023), we evaluated motion quality using embeddings from a classifier to compute several metrics: Fréchet Inception Distance (FID), Diversity, Multimodality, Intent Recognition Accuracy (IRA). Additionally, we report Average Variance Error (AVE), which quantifies the discrepancy in joint position variance relative to the ground truth. We trained the classifier to predict interaction intent based on full-body joint positions, including hand joints, and object trajectories. See more details in the Appendix.
Regarding the grasp accuracy itself, we identify two primary failure modes: object-scene/human penetration and support-less object floating. To quantify Penetration, we calculated the mean penetration depth in mm, considering only frames where penetration occurred. For Floating, we employed a similar method, but only considered frames without penetration.
In all our evaluation we considered clips of the motions in which contact occurs, re-sampled to 30 frames.
5.2. Quantitative Results
Table 1 compares our method to IMoS. HOIDiNi achieves better FID and IRA, indicating improved realism, and significantly reduces floating while maintaining comparable penetration levels.
User Study. We conducted a user study with 24 participants, evaluating, in total, 12 side-by-side randomly selected samples of two models using the same inputs. As shown in Figure 7, users preferred the results generated by our framework. A representative screenshot from the study interface is shown in Figure 10.
Figure 7. User Study. We compare to two baselines, measuring on grasp quality, prompt adherence, and overall quality over 12 random samples, each evaluated by at least 10 users. The dashed line marks 50% ratio.
The effectiveness of the two-phase design is demonstrated in Figure 8, when compared to the single-phase setup, in which contact predictions are optimized jointly with human motion. This single-phase ablation results in frequent updates to contact assignments throughout the process. These updates shift the contact loss objective over time, making the optimization less stable and harder to converge. In contrast, our two-phase approach separates contact prediction and motion optimization: contact pairs are predicted and fixed in the first (object-centric) phase, allowing the second (human-centric) phase to optimize a stable and well-defined contact loss objective. This decoupling enables more consistent optimization behavior and improves convergence, as reflected in Table 1.
5.3. Qualitative Results
The supplementary video showcases a variety of motions generated by our model, along with visual comparisons to baseline methods. Figure 5 shows contact pairs generated by HOIDiNi. Figure 6 presents an interaction generated by HOIDiNi for the prompt “The person is playing with the gaming controller.” The IMoS baseline demonstrates a common failure where contacts are lacking semantic meaning and physical plausibility due to the fixed-contact snapping approach. Similarly, only replacing our predicted contact pairs with nearest-neighbor assignments also results in incorrect contacts and implausible motion. We further witness that applying our losses through classifier guidance improves hand-object proximity compared to inference-only, but does not produce realistic interactions. Figure 4 presents additional examples generated by our method. These results further highlight the semantic correctness and physical plausibility of the synthesized motions across a diverse set of prompts and object types.
Figure 8. Single vs. Two-Phase Optimization. Comparison of contact loss and changes in predicted contact bits during DNO optimization (shared y-axis). (Left): In the single-phase setup, contact predictions evolve alongside motion, causing unstable objectives and hindering convergence. (Right): In our two-phase approach, contact pairs are fixed after the object-centric phase, resulting in stable contact loss during the human-centric phase presented here. In this example, the two-phase setting converges to a 10× lower value. Contact-Bits Changes denote the number of bit flips in the predicted contact matrix between successive steps.
6. Discussion
This paper presented a method for diffusion-based HOI generation. The main challenge with HOI generation is realistic and accurate contacts. We observe explicit optimization to satisfy this demand tends to drift the generation away from the learned manifold, yielding implausible results. At the same time, classifier guidance remains closer to the manifold, but fails to adhere to precise contacts. To address this, we introduce DNO in the context of HOI for the first time. We also observe that heuristically determining the contact pairs (e.g, through nearest neighbor projection) is also a crucial pitfall for realism and plausibility. Predicting these pairs, however, turned out to be infeasibly challenging due to convergence issues. This drove the design of a variable-separated optimization scheme, consisting of the two phases. As attested through our user study, our approach represents a step forward in HOI generation, balancing realistic movement, accurate contacts, and superior human likeness.
We believe this exciting avenue opens up many interesting questions for future work. Firstly, all our experiments are confined to the GRAB dataset, to emphasize superiority compared to current art. New and larger datasets are recently emerging, and we encourage future investigation to incorporate them in HOI modeling, including general, not HOI related, human motion data. Similarly, our method is confined to the handful of objects seen during training. Expanding to novel shapes, perhaps using very recent shape foundational models (Xiang et al., 2024), seems equally promising. Lastly, like previous methods our approach is optimization-based, which renders its generation long, and applicability limited. A pure sampling-based method would greatly benefit the community. The aforementioned larger datasets could facilitate this goal as well.
No matter the direction, we argue future investigation should not compromise surface contact accuracy nor human likeness, and believe our Contact-Pairs-based motion outline approach will continue to prove beneficial in future developments.
Acknowledgements.
This research was partly supported by the Israel Science Foundation (1337/22), Len Blavatnik, and the Blavatnik family foundation.
Appendix
Appendix A Contact Representation
Directly generating index pairs via diffusion is challenging. To address this, we adopt a more learnable representation. We define a fixed anchor set $\mathcal{A}$, consisting of a subset of MANO (Romero et al., 2017a) hand vertices located on the palm (Figure 9). At each frame $f$, and for each anchor $a\in\mathcal{A}$, a binary variable $b_{a}$ indicates whether the anchor is in contact. A corresponding position $p_{a}\in\mathbb{R}^{3}$ specifies the contact location on the object surface in its rest pose. This yields the contact representation:
Appendix B Evaluation Metrics

### v2

@@ 4.2 CPHOI Diffusion Model
CPHOI generates the next motion segment $s^{n}=[F_{i}]_{i=1}^{L}$ of $L$ frames, given the previous segment, $s^{n-1}$, the object geometry in point-cloud format, and a text prompt that describes the interaction. For each denoising step $t\in[0,T]$, the inputs to the transformer decoder backbone are the frames of the previous segment $s^{n-1}$, followed by the current segment to be denoised, $s^{n}_{t}$. The model predicts the clean version of the segment, $\hat{s}^{n}_{0}$.
The object's geometry, the text condition, and the timestep $t$ are encoded into a sequence of latent embeddings and injected through the cross-attention of each transformer layer, at each denoising step. To encode the geometry of the interacted object, we use a shape encoder with a PointNet++ (Qi et al., 2017) architecture, which is trained simultaneously with the diffusion model. We uniformly sample the mesh using $V$ points, and encode each one into a latent descriptor of length $C$ ($V=512$ and $C=512$ in all our experiments). The text prompt is encoded into a sequence of embeddings using a pre-trained and fixed DistilBERT (Sanh, 2019) model. The diffusion timestep $t$ is embedded using a standard positional embedding denoted $E_{t}$.
@@
4.3 Two-phase Generation Optimization
Accurate body-object contacts are a key component of plausible HOI motion, but represent a discontinuous space, with discrete and unstable decision making. Thus we found the straightforward optimization, using a single step challenging in practice (see Figure 8). To address this, we separate the process into two phases, each solving a different part of the problem variables:
Phase 1: Object-Centric. In this phase, we optimize the object related part only to outline the motion, based on the given prompt, object geometry, and scene constraints (e.g., table surface). Although the full output of the model is the triplet (CP,H,O)(CP,H,O)—representing the contact-pair sequence, human motion, and object motion respectively—only $(CP)$ and $(O)$ are considered in this stage. We found weight sharing between phases to improve performance. The objective $\mathcal{L}_{Object}$ for this phase is defined to avoid object-scene penetrations and optionally to guide the object toward a set of target poses in specified frames (as sometimes used in the literature  (Li et al., 2024a)):
@@ 4.3 Two-phase Generation Optimization
This objective comprises a penetration term between the object and the scene $\mathcal{L}_{P_{OS}}$, a goal term encouraging the object to achieve specified poses in specified times $\mathcal{L}_{Goal}$, and the static term, $\mathcal{L}_{Static}$, that encourage the object to stay static in frames without predicted contact. The last two are detailed in Appendix C.
Penetration Loss. To measure interpenetration between two watertight meshes, denoted $M_{a}$ and $M_{b}$, we define a bidirectional penetration loss that penalizes vertices of one mesh that are located inside the volume enclosed by the other. Each mesh $M$ consists of a set of vertices $\mathcal{V}$ and a set of polygons $\mathcal{F}$. Each polygon $f\in\mathcal{F}$ is associated with an outward-facing normal vector $\mathbf{n}_{f}$.
Figure 6: Comparisons. HOIDiNi generates semantically correct and accurate interaction with the gaming controller. From left to right: IMoS (Ghosh et al., 2023) optimization yields inferior contacts and unrealistic motion; CPHOI inference only generates decent poses but fails to satisfy contacts; Our losses with Classifier Guidance brings the object insufficiently closer. Replacing our contact-point prediction with the popular nearest-neighbor heuristic fails to choose correct contacts, in contrast to our plausible and human-like result.
For each vertex $\mathbf{v}\in\mathcal{V}_{a}$, we identify whether it is inside the other mesh simply through randomly casting a single ray, and checking the normal direction of the intersected triangle if it exists. We denote the vertices found as inside $\mathcal{V}_{a}^{\text{in}}\subset\mathcal{V}_{a}$.
@@
Hence, the full symmetric loss is, $\mathcal{L}_{\text{Penetration}}=\mathcal{L}_{\text{pen}}(M_{a}\rightarrow M_{b})+\mathcal{L}_{\text{pen}}(M_{b}\rightarrow M_{a})$.
Phase 2: Human-Centric Phase. This phase refines the human motion to conform to the fixed object motion and contact sequence $(CP,O)$ produced in Phase 1. With the contact points fixed, this optimization over the human manifold is much more stable. The objective of this diffusion-based optimization is defined as:
representing human-object, human-scene, and human-human penetration losses, respectively. Specifically, $\mathcal{L}_{P_{HO}}$ penalizes interpenetration between the human and the object, $\mathcal{L}_{P_{HS}}$ prevents collisions with the static scene, and $\mathcal{L}_{P_{HH}}$ reduces self-intersections within the human mesh, particularly between the hands. The contact loss $\mathcal{L}_{C}$ uses $L_{2}$ distance to push hand vertex locations to targets on the moving object, as dictated by $CP$ and $O$ from the previous phase. The penetration loss terms are similar to the penetration term $\mathcal{L}_{P_{OS}}$ defined in phase 1. Together, these losses represent physical plausibility, encouraging surface contacts while avoiding penetration, and the optimization scheme ensures realism and human likeness.
GT 0.995 74.6% 0.194 5.0 $\pm$ 1.8 2.6 $\pm$ 1.8
IMoS 0.205 1.026 0.121 46.5% 0.204 3.0 $\pm$ 7.7 52.2 $\pm$ 53.8
HOIDiNi (Ours) 0.159 0.996 0.121 62.3% 0.245 6.8 $\pm$ 2.9 2.3 $\pm$ 4.7
Single-Phase 0.221 1.026 0.140 43.0% 0.241 7.8 $\pm$ 3.6 20.1 $\pm$ 21.7
Higher Penetration Coef. $\lambda_{P_{HO}}$ 0.156 1.015 0.144 57.0% 0.225 5.6 $\pm$ 2.9 5.9 $\pm$ 7.2
NN instead of Contact Pairs 0.162 1.040 0.153 57.9% 0.245 7.7 $\pm$ 3.8 4.3 $\pm$ 5.0
No Jitter Loss 0.160 1.026 0.143 60.5% 0.220 7.1 $\pm$ 2.7 1.9 $\pm$ 4.0
Table 1: Quantitative Results and Ablation Study. Comparison on the GRAB dataset (Taheri et al., 2020b) using IMoS-defined (Ghosh et al., 2023) metrics for motion realism, along with average floating and penetration errors (in millimeters). $\rightarrow$ indicates that better is closer to ground-truth performance.
5 Experiments
Figure 7: User Study. We compare to two baselines, measuring on grasp quality, prompt adherence, and overall quality over 12 random samples, each evaluated by at least 10 users. The dashed line marks 50% ratio.
5.1 Evaluation Setting
Implementation Details.
CPHOI is implemented as a Transformer decoder architecture with 8 layers and a hidden dimension of 512. Our point-wise object embedding network is a PointNet++  (Qi et al., 2017) fed with 512 randomly sampled vertices from the conditioned object. We condition the motion on a prefix of 15 frames and generate 100 frames. The model is trained with DDPM (Ho et al., 2020b); DDIM (Song et al., 2020) is used at inference. Further details are at Appendix A.
Data.
We evaluate our approach using the GRAB (Taheri et al., 2020a) and OMOMO (Li et al., 2023) datasets. GRAB includes SMPL-X (Pavlakos et al., 2019a) human motion parameters, object 6DoF trajectories, and per-vertex contact force data for both human and object meshes. The dataset comprises 1,334 motion samples from 10 different subjects. Since different human subjects in the dataset exhibit varying shape parameters $\beta$, and considering the dataset’s relatively small scale compared to datasets like HumanML3D, we re-target all motions to a standardized, neutral SMPL-X model with shape parameters fixed to $\beta=0$. Additionally, we used ChatGPT 4o to enhance its discrete action annotations into text prompts. The OMOMO dataset includes 27,952 motion sequences, interacting with 13 objects. It does not include finger motions and hence doesn't enable precise interactions. We use it strictly for comparisons with the recent literature  (Li et al., 2024a). Following it, we use the data's subset, called FullBodyManipulation.
Baselines
We compare our approach against IMoS (Ghosh et al., 2023), CHOIS (Li et al., 2024a), and several internal baselines. One baseline applies our model at inference time without any optimization. Another replaces the diffusion noise optimization with classifier guidance. The third substitutes the predicted contact pairs with nearest-neighbor-based pairs. To enable a fair comparison with classifier guidance, we adopt DDPM hyper-sampling and increase the number of denoising steps from 8 to 500, consistent with Peng et al. (Peng et al., 2023). We report results using the optimal guidance scale. For the nearest-neighbor baseline, contact assignment is updated at each optimization step based on the current nearest neighbors. Additionally, we explore increasing the penetration loss weight by setting $\lambda_{P_{HO}}=0.9$.
Metrics. We evaluated our method on motion realism and interaction accuracy. For GRAB, we followed IMoS (Ghosh et al., 2023) metrics, computing FID, Diversity, Multimodality, Intent Recognition Accuracy (IRA), and Average Variance Error (AVE) using a classifier trained on full-body joint positions, hand joints, and object trajectories (Appendix D). Grasp accuracy was measured via two failure modes: Penetration (mean depth in mm, frames with penetration only) and Floating (frames without penetration). For OMOMO, we used the CHOIS benchmark, covering condition adherence, motion fidelity, and interaction accuracy (Appendix D).
Method Condition Matching Human Motion Interaction
$T_{s}\downarrow$ $T_{e}\downarrow$ $T_{xy}\downarrow$ $H_{\text{feet}}\downarrow$ FS $\downarrow$ $R_{\text{prec}}\uparrow$ FID $\downarrow$ $C_{\text{prec}}\uparrow$ $C_{\text{rec}}\uparrow$ $_{F_{1}}\uparrow CC_{\%}$ $P_{\text{hand}}\downarrow$
CHOIS 1.90 6.90 2.81 4.48 0.34 0.43 0.97 0.80 0.64 0.67 0.56 0.61
HOIDiNi (ours) 0.00 0.00 0.00 3.17 0.30 0.42 1.24 0.78 0.8 0.77 0.76 0.67
Table 2: Comparison to CHOIS (Li et al., 2024a) over the OMOMO (Li et al., 2023) dataset. Measuring condition matching, human motion, and interaction via a set of metrics defined by CHOIS.
Figure 8: Single vs. Two-Phase Optimization. Comparison of contact loss and changes in predicted contact bits during DNO optimization (shared y-axis). (Left): In the single-phase setup, contact predictions evolve alongside motion, causing unstable objectives and hindering convergence. (Right): In our two-phase approach, contact pairs are fixed after the object-centric phase, resulting in stable contact loss during the human-centric phase presented here. In this example, the two-phase setting converges to a 10× lower value. Contact-Bits Changes denote the number of bit flips in the predicted contact matrix between successive steps.
5.2 Results
Table 1 compares our method to IMoS. HOIDiNi achieves better FID and IRA, indicating improved realism, and significantly reduces floating while maintaining comparable penetration levels. Table 2 compare our model to CHOIS using the metrics defined by them and shows that HOIDiNi improves the contact accuracy while maintaining comparable motion fidelity. The condition matching metrics demonstrate the preciseness of the DNO mechanism, which consistently delivers zero error.
User Study. We conducted a user study for the GRAB dataset with 24 participants, evaluating 12 side-by-side randomly selected samples of two models using the same inputs. As shown in Figure 7, users preferred the results generated by our framework. A representative screenshot from the study interface is shown in Figure 10.
The two-phase design is investigated in Figure 8, comparing to the single-phase setup, where contact predictions are optimized jointly with human motion. The single-phase approach results in frequent updates to contact assignments throughout the process. These updates shift the contact loss objective over time, making the optimization less stable and harder to converge. In contrast, our two-phase approach separates contact prediction and motion optimization: contact pairs are predicted and fixed in the first (object-centric) phase, allowing the second (human-centric) phase to optimize a stable and well-defined contact loss objective. This decoupling enables more consistent optimization behavior and improves convergence, as reflected in Table 1.
Qualitative Results. The supplementary video showcases a variety of motions generated by our model, along with visual comparisons to baseline methods. Figure 3 shows contact pairs generated by HOIDiNi. Figure 6 presents an interaction generated by HOIDiNi for the prompt “The person is playing with the gaming controller.” The IMoS baseline demonstrates a common failure where contacts are lacking semantic meaning and visually plausibility due to the fixed-contact snapping approach. Similarly, only replacing our predicted contact pairs with nearest-neighbor assignments also results in incorrect contacts and implausible motion. We further witness that applying our losses through classifier guidance improves hand-object proximity compared to inference-only, but does not produce realistic interactions. Figure 5 presents additional examples generated by our method. These results further highlight the semantic correctness and visually plausibility of the synthesized motions across a diverse set of prompts and object types.
6 Conclusions
We introduced HOIDiNi, an approach for high-precision Human–object Interaction that provides motion fidelity through Diffusion Noise Optimization. Our results demonstrate accurate contact handling and natural motion across complex interaction scenarios. Beyond HOI, we view HOIDiNi as a platform for cases where high-precision is required and regular diffusion inference fails. We encourage the community to use HOIDiNi to advance controllable, high-fidelity motion generation.
Appendix
Appendix A Implementation Details
CPHOI is implemented as a Transformer decoder architecture with 8 layers and a hidden dimension of 512. Our point-wise object embedding network is a PointNet++  (Qi et al., 2017) fed with 512 randomly sampled vertices from the conditioned object. The model is trained with DDPM (Ho et al., 2020b); DDIM (Song et al., 2020) is used at inference. More details per experiment can be found in Table 3. We use the Adam optimizer (Kingma, 2014) for the noise optimization procedure. To reduce memory costs, we focus only on the MANO  (Romero et al., 2017a) subset of vertices of the SMPL-X human body and simplify it further to 1,100 vertices for each hand. The object meshes are simplified to 3,000 faces.
Appendix B Contact Representation
Directly generating index pairs via diffusion is challenging. To address this, we adopt a more learnable representation. We define a fixed anchor set $\mathcal{A}$, consisting of a subset of MANO (Romero et al., 2017b) hand vertices located on the palm (Figure 10). At each frame $f$, and for each anchor $a\in\mathcal{A}$, a binary variable $b_{a}$ indicates whether the anchor is in contact. A corresponding position $p_{a}\in\mathbb{R}^{3}$ specifies the contact location on the object surface in its rest pose. This yields the contact representation:
For the OMOMO (Li et al., 2023) dataset, which lacks fingers' motion, we follow CHOIS and define the middle finger in each hand in the SMPL-X body model as the anchor, resulting in only two anchors for this benchmark.
Experiment GRAB [2020a] OMOMO [2023]
Training parameters #\# input prefix frames 15 1
#\# generated frames 100 119
diffusion steps ($T$) 8 14
training steps 120K 50K
batch size 64 64
DNO parameters perturbation scale $10^{-6}$ $10^{-5}$
difference penalty $10^{-6}$ $10^{-5}$
$\lambda_{C}$ 0.95 0.95
$\lambda_{Foot}$ 0.5 0.5
$\lambda_{Jitter}$ $10^{-5}$ $10^{-3}$
$\lambda_{P_{HO}}$ 0.05 0.05
$\lambda_{P_{HS}}$ 0.2
$\lambda_{P_{HH}}$ 0.05 0.05
$\lambda_{P_{OS}}$ 1.2 0.05
$\lambda_{Goal}$ 0.5 0.9
$\lambda_{Static}$ 0.9 0.05
$\lambda_{FeetFloorContact}$ 0.5
Table 3: Hyper-parameters in use for each experiment.
Appendix C Additional DNO Loses
C.1 Object-centric losses
@@
C.2 Human-centric losses
Feet-floor Contact Loss. For the OMOMO experiment, following CHOIS, we add a loss term that enforces accurate foot contact at the mesh level.
When reconstructing the human mesh with SMPL-X (Pavlakos et al., 2019a) using predicted root positions, joint rotations, and subject-specific body parameters, the generated feet may occasionally fail to touch the floor. To address this, we add a guidance term that encourages realistic feet-floor contact.
Let $\bm{J}_{l}$ and $\bm{J}_{r}$ denote the positions of the left and right toe joints. At each frame, the supporting foot is identified by comparing their z-coordinates. We further set a threshold height $h=0.02$ meters, derived from analyzing foot heights in the ground truth motion. The guidance term is then defined as:
$L_{\text{FeetFloorContact}}=|\min(\bm{J}_{l}^{z},\bm{J}_{r}^{z})-h|_{2}.$ (1)
This measures the vertical deviation between the lower toe joint and the threshold height $h$.
Appendix D Evaluation Metrics
For both benchmarks, GRAB and OMOMO, we evaluate our method along two dimensions: motion realism and interaction accuracy.
D.1 GRAB Evaluation